# Rethinking LLM Memorization through the Lens of Adversarial Compression

### v1

@@ 3.2 MiniPrompt: A Practical Test for Compressible Memorization
Since the compression rate $ACR$ is defined in terms of the solution to a minimization problem, we propose an approximate solver to estimate compression rates, see Algorithm 1. Specifically, to find the minimal prompt for a particular target sequence, or to solve the optimization problem in Equation 2, we use GCG [^1] The full GCG algorithm is outlined in Appendix A. (Zou et al., 2023) and search over sequence length. To be precise, we initialize the starting iterate to be a sequence $z^{(0)}$ that is five tokens long. Each step of our algorithm runs GCG to optimize $z$ for $n$ steps. If the resulting prompt successfully produces the target string, i.e. $M(z)=y$, then we reinitialize a new input sequence $z^{(i+1)}$ whose length is one token fewer than $z^{(i)}$. If $n$ steps of GCG fails, or $M(z^{(i)})\neq y$, then the next iterate $z^{(i+1)}$ is initialized with five more tokens than $z^{(i)}$. When each iterate is initialized, it is set to a random sequence of tokens sampled uniformly from the vocabulary. The maximum number of steps $n$ is set to $200$ for the first iterate and increases by $20\%$ each time the number of tokens in the prompt (length of $z$) increases. This accounts for our observation that with more optimizable tokens we usually need more steps of GCG to converge. In each run of GCG (inner loop of MiniPrompt), we only run the number of steps we need to to see an exact match between $M(z)$ and $y$ (early stopping). Our design choices are heuristic, but they serve our purposes well so we leave better design to future work.
@@ 4 Compressible Memorization in Practice
We show the practical value of our definition and algorithm through several case studies and we show that the definition meets our expectations around memorization with validation experiments. Our case studies start with a demonstration of how a model owner trying to circumvent a regulation about data memorization might use in-context unlearning (Pawelczyk et al., 2023), or design-specific system prompts that change how apparent is memorization. Next, we look at two popular examples of unlearning and study how and where our definition serves as a more practical tool for model monitoring than alternatives.
@@
Рис. 3: In-Context Unlearning (ICUL) fools completion not compression. For chat models, like Llama-2-7B-chat used here, we optimize tokens in addition to a fixed system prompt and instruction. In this setting, we show that MiniPrompt compresses the quote in purple to the two blue tokens in the prompt in the top cell. Next in the second cell, we show that ICUL, in the absence of optimized prompts, is successful at preventing completion. Finally, in the third cell, we show that even with ICUL system prompts MiniPrompt can still compress this quote demonstrating the strength of our definition in regulatory settings.
@@ 4.1 The Illusion of Compliance
As data usage regulation advances, there is an emerging motive for organizations and individuals that serve or release models to make it hard to determine that their models have memorized anything. To that end, we consider a simple defense that these model owners might employ as a proof-of-concept that one can easily fool existing definitions of memorization but not our compression-based definition. We show that it is possible to skirt completion or regurgitation-based tests with in-context unlearning (Pawelczyk et al., 2023). Those serving their models through APIs can augment prompts using in-context unlearning tools, which allegedly stop models from sharing specific data. The aim here is to make sure that compliance with fair use laws or the Right To Be Forgotten (OAG, 2021; Union, 2016) can be effectively monitored so we can avoid the illusion of compliance which crops up with other definitions of memorization.
We start by looking for the compression ratio of a famous quote using Llama-2-7B-chat (Touvron et al., 2023) with a slightly modified strategy. Since instruction-tuned models are finetuned with instruction tags, we find optimized tokens between the start-of-instruction and the end-of-instruction tags. Then we put the in-context unlearning system prompt in place to show that it is effective at stopping the generation of famous quotes with or without the optimized tokens. Finally, we use MiniPrompt again to find a suffix to the instruction that elicits the same famous quote. In Figure 3, we show examples of each of these steps.
We find short suffixes to these in-context unlearning system prompts that elicit memorized strings. Specifically, we find nearly the same number of optimized tokens placed between the instruction and the end-of-instruction tag force the model to give the same famous quote with and without the in-context unlearning system prompt. This consistency in the compression ratio—and therefore the memorization test—matches our intuition that without changing model weights memorized samples are not forgotten. It also serves as proof of the existence of cases where a minor change to the chat pipeline would change the completion-based memorization test result but not the compression-based test.
@@ 4.2 TOFU: Unlearning and Memorization with Author Profiles
As unlearning progresses, we prompt the model to generate answers to the supposedly unlearned questions and record the portion of data that can be completed and compressed. Figure 4 shows that after only 16 unlearning steps, none of the unlearned questions can be completed exactly. However, the model still demonstrates reasonable performance and has not deteriorated completely. As expected, compression shows that a considerable amount of the ‘forget data’ is compressible and hence memorized. This case suggests that we cannot safely rely on completion as a metric for memorization because it is too strict.
@@
Рис. 4: Left: Completion vs compression on TOFU data, unlearning Phi-1.5 with gradient ascent. Right: Generation after 20 unlearning steps.
@@ 4.3 Trying to Forget Harry Potter
In their paper on unlearning Harry Potter, Eldan & Russinovich (2023) claim that Llama-2-chat can forget about Harry Potter with several steps of unlearning. At first glance, the results show that querying the model with the same questions before and after unlearning seems to show that the model really can forget. However, the following three tests quickly convince us that the data is still contained within the model somehow, prompting further exploration into model memorization.
@@
    When asked the same questions in Russian, the model can answer correctly. We provide examples of such behavior in Appendix C.
@@
    With adversarial attacks designed to force answers in the affirmative without any information about the true answer, we can elicit the correct response, see Figure 6.
@@
Рис. 5: Negative log-likelihood (normalized to $[0,1]$) of true and false answers given a Harry Potter question. Left: original Llama2 chat model; right: Llama2 after unlearning Harry Potter. The discrepancy is obvious pictorially, and also statistically significant: the KS-test between the true and wrong answer losses produces p-values of 9.7e-24 and 5.9e-14, respectively.
[…]
Рис. 6: Left: Fraction of Harry Potter texts that are compressible. Right: an example of hard tokens that elicit Harry Potter text.
@@ 4.4 Bigger Models Memorize More
Since prior work has proposed alternative definitions of memorization that show that bigger models memorize more (Carlini et al., 2023), we ask whether our definition leads to the same finding. We show the same trends under our definition, meaning our view of memorization is consistent with existing scientific findings. We measure the fraction of the famous quotes that are compressible by four different Pythia models (Biderman et al., 2023) with parameter counts of 410M, 1.4B, 6.9B, and 12B and the results are in Figure 7).
@@ 4.5 Validation of MiniPrompt with Four Categories of Data
We look at random sequences of tokens because we want to rule out the possibility that we can always find an adversarial, few-token prompt even for random output—random strings should not be compressible. To this end, we draw uniform random samples with replacements from the token vocabulary to build a set of 100 random outputs that vary in length (between 3 and 17 tokens). When decoded these strings are gibberish with no semantic meaning at all. We find that these strings are never compressible—that is across multiple model sizes we never find a prompt shorter than the target that elicits the target sequence as output, see the zero-height bars in Figure 8.
Рис. 7: Memorization in Pythia models. Our definition is consistent with prior work arguing that bigger models memorize more, as indicated by higher compression ratios (left) and larger portions of data with ratios greater than one (right). These figures are from the Famous Quotes dataset.
Рис. 8: Memorization in Pythia-1.4B. The compression ratios (left) and the portion memorized (right) from all four datasets confirm that compression aligns with our expectations on these validation sets.
@@
To further determine the validity of our definition, we investigate the average compression rate of natural text that is not in the training set. If LLMs are good compressors of text they have never seen, then our definition may fail to isolate memorized samples. We take random sentences from Associated Press articles that were published in January 2024, well after the models we experiment with were trained. These strings are samples from the distribution of training data as the training set includes real news articles from just a few months prior. Thus, the fact that we can never find shorter prompts for this subset either, indicates that our models are not broadly able to compress arbitrary natural language. Again, see the zero-height bars in Figure 8.
@@
Next, we turn our attention to famous strings, of which many should be categorized as ‘memorized’ by any useful definition. These are quotes like ‘‘to be, or not to be, that is the question,’’ which we know are examples that are repeated many times in the training data. We find that Pythia-1.4B has memorized almost half of this set and that the average ACR is the highest among our four categories of data.
@@
Finally, as our first experimental inquiry, we look at the memorization of training samples that are not common or famous, but that do exist in the training set. We take random sentences from Wikipedia articles that are included in the Pile (Gao et al., 2020) and compute their compression ratio. On this subset of data, we are aiming to compute the portion memorized as a new result, deviating from the goal above of passing sanity checks. Figure 8 shows that some of these sentences from Wikipedia are memorized and that the average compression ratio is between the average among famous quotes and news articles. Note that the memorized samples form this subset are strings that appear many times on the internet like ‘‘The Burton is a historic apartment building located at Indianapolis, Indiana.’’
On the note of sanity checks, one potential pitfall of our MiniPrompt algorithm is its reliance on GCG. It is possible that there exist shorter strings than we can find. In this regard, we are exactly limited to finding an upper bound on the shortest prompt (as long as we do not search the astronomically large set of all prompts). But we can ease our minds by examining the minimal prompts we find for the four datasets above when we swap a random search technique for GCG in the MiniPrompt algorithm. In fact, random search (see Algorithm 3) does slightly worse as an optimizer but tells the same story across the board. Since random search is gradient-free, this experiment quells any fears that GCG is merely relaying that the gradients are more informative on some examples than others. The details of this experiment and our exact random search algorithm are in Appendix A.
5 Additional Related Work
Prompt Optimization
We borrow prompt optimization tools from work on jailbreaking where the goal is to force LLMs to break their alignment and produce nefarious and toxic output by way of optimizing prompts (Zou et al., 2023; Zhu et al., 2023; Chao et al., 2023; Andriushchenko, 2023). Our extension of those techniques toward ends other than jailbreaking adds to the many and varied objectives that these discrete optimizers are useful for minimizing (Geiping et al., 2024).
Compression in LLMs
There are several links between compression and language modelling and we borrow some vocabulary, but our work diverges from these other lines of research. For example, Delétang et al. (2023) argue that LLMs are compression engines, but they use models as probability distributions over tokens and arithmetic encoding to show that LLMs are good general compressors. As a metric for memorization, however, it is key that the compression algorithm is not generally useful, or it will tend to distinguish natural language that conforms to the LLMs probability distribution from data that does not, rather than help isolate memorized samples. Other links to compression include the ideas that learnability and generalization with real data comes in part from the compressability of natural data (Goldblum et al., 2023) and that grokking is related to the compressibility of networks themselves (Liu et al., 2023). Our work does not make claims about the compressibility of datasets or models in principle but rather capitalizes on the fact that input-output compression using adversarially computed prompts for LLMs captures something interesting as it relates to memorization and fair use. In fact, Jiang et al. (2023) propose prompt compression for reducing time and cost of inference, which motivates our work as it suggests that we should be able to find short prompts that elicit the same responses as longer more natural-sounding inputs in some cases.
Unlearning
The focus of machine unlearning (Bourtoule et al., 2021; Sekhari et al., 2021; Ullah et al., 2021) is to remove private, sensitive, or false data from models without retraining them from scratch. Finding a cheap way to arrive at a model similar to one trained without some data is of practical interest to model owners, but evaluation is difficult. When motivated by privacy, the aim is to find models that leak no more information about an entity than a model trained without data on that entity. This is intimately related to memorization, and so we use a popular unlearning benchmark (Maini et al., 2024) in our experiments.
@@ 6 Discussion
When proposing new definitions, we are tasked with justifying why a new one is needed and as well as showing its ability to capture a phenomenon of interest. This stands in contrast to developing detection/classification tools whose accuracy can easily be measured using labeled data. It is a difficult task by nature to define memorization as there is no set of ground truth labels that indicate which samples are memorized. Consequently, the criteria for a memorization definition should rely on how useful it is. Our definition is a promising direction for future regulation on LLM fair use of data as well as helping model owners confidently release models trained on sensitive data without releasing that data.
From an intuitive perspective, one can imagine compression as a communication game—Alice tries to send some data to Bob with as few bits as possible. In our regime, we assume that both Alice and Bob have access to the same LLM $M$ as the compressor. If Alice wants to send some string $y$ to Bob, she can instead send the minimal prompt $x$ and Bob can decode $y=M(x)$. When $|x|\leq|y|$, Alice has successfully compressed the data $y$. Future work to explore other ways of measuring compression and faster/cheaper optimization approaches could be fruitful. One drawback of optimizing hard tokens is the intrinsic complexity of discrete optimization. Instead, if we optimize over the continuous soft token space, the process will be much more efficient, but it remains unclear how to measure the information content of a soft token.
@@
  • Biderman et al. (2023) Stella Biderman, Hailey Schoelkopf, Quentin Gregory Anthony, Herbie Bradley, Kyle O’Brien, Eric Hallahan, Mohammad Aflah Khan, Shivanshu Purohit, USVSN Sai Prashanth, Edward Raff, et al. Pythia: A suite for analyzing large language models across training and scaling. In International Conference on Machine Learning, pp.  2397–2430. PMLR, 2023.
  • Bourtoule et al. (2021) Lucas Bourtoule, Varun Chandrasekaran, Christopher A Choquette-Choo, Hengrui Jia, Adelin Travers, Baiwu Zhang, David Lie, and Nicolas Papernot. Machine unlearning. In 2021 IEEE Symposium on Security and Privacy (SP), pp.  141–159. IEEE, 2021.
  • Carlini et al. (2019) Nicholas Carlini, Chang Liu, Úlfar Erlingsson, Jernej Kos, and Dawn Song. The secret sharer: Evaluating and testing unintended memorization in neural networks. In 28th USENIX security symposium (USENIX security 19), pp.  267–284, 2019.
@@
  • Eldan & Russinovich (2023) Ronen Eldan and Mark Russinovich. Who’s harry potter? approximate unlearning in llms. arXiv preprint arXiv:2310.02238, 2023.
@@
  • Ippolito et al. (2023) Daphne Ippolito, Florian Tramèr, Milad Nasr, Chiyuan Zhang, Matthew Jagielski, Katherine Lee, Christopher A Choquette-Choo, and Nicholas Carlini. Preventing generation of verbatim memorization in language models gives a false sense of privacy. In Proceedings of the 16th International Natural Language Generation Conference, pp.  28–53. Association for Computational Linguistics, 2023.
@@
  • (21) Cade Metz and Katie Robertson. Openai seeks to dismiss parts of the new york times’s lawsuit. The New York Times. URL https://www.nytimes.com/2024/02/27/technology/openai-new-york-times-lawsuit.html#:~:text=In%20its%20suit%2C%20The%20Times,someone%20to%20hack%20their%20chatbot.
@@
  • Shokri et al. (2017) Reza Shokri, Marco Stronati, Congzheng Song, and Vitaly Shmatikov. Membership inference attacks against machine learning models. In 2017 IEEE symposium on security and privacy (SP), pp.  3–18. IEEE, 2017.
@@
  • Ullah et al. (2021) Enayat Ullah, Tung Mai, Anup Rao, Ryan A Rossi, and Raman Arora. Machine unlearning via algorithmic stability. In Conference on Learning Theory, pp.  4126–4142. PMLR, 2021.
@@
Приложение A Algorithms In Our Experiments
@@
Algorithm 2 Greedy Coordinate Gradient (GCG) (Zou et al., 2023)
@@
Algorithm 3 Random Search (for LLM prompts) (Andriushchenko, 2023)
@@
Рис. 9: Pythia-1.4B Memorization with Random Search.
Рис. 10: Pythia-410M Memorization with GCG.
Приложение B More Details of In-context Unlearning
@@
Приложение C More Details of Unlearning Harry Potter

### v2

@@ 3.2 MiniPrompt: A Practical Algorithm for Compressible Memorization
Since the compression rate ACR is defined in terms of the solution to a minimization problem, we propose an approximate solver to estimate compression rates, see Algorithm 1. Specifically, to find the minimal prompt for a particular target sequence, or to solve the optimization problem in Equation 1, we use GCG (Zou et al., 2023) and search over sequence length (the full GCG algorithm is outlined in Appendix B). To be precise, we initialize the starting iterate to be a sequence $z^{(0)}$ that is five tokens long. Each step of our algorithm runs GCG to optimize $z$ for $n$ steps. If the resulting prompt successfully produces the target string, i.e. $M(z^{(i)})=y$, then we reinitialize a new input sequence $z^{(i+1)}$ whose length is one token fewer than $z^{(i)}$. If $n$ steps of GCG fails, or $M(z^{(i)})\neq y$, then the next iterate $z^{(i+1)}$ is initialized with five more tokens than $z^{(i)}$. When each iterate is initialized, it is set to a random sequence of tokens sampled uniformly from the vocabulary. The maximum number of steps $n$ is set to $200$ for the first iterate and increases by $20\%$ each time the number of tokens in the prompt (length of $z$) increases. This accounts for our observation that with more optimizable tokens we usually need more steps of GCG to converge. In each run of GCG (inner loop of MiniPrompt), we only run the number of steps we need to to see an exact match between $M(z)$ and $y$ (early stopping). Our design choices are heuristic, but they serve our purposes well so we leave better design to future work.
@@ 4 Compressible Memorization in Practice
We show the practical value of our definition and algorithm through several case studies as well as that the definition meets our expectations around memorization with validation experiments. Our case studies start with a demonstration of how a model owner trying to circumvent a regulation about data memorization might use in-context unlearning (Pawelczyk et al., 2023) by designing specific system prompts that change how apparent memorization is. Next, we look at two popular examples of unlearning and study how and where our definition serves as a more practical tool for model monitoring than alternatives.
@@
Рис. 2: In-Context Unlearning (ICUL) fools completion not compression. For chat models, like Llama-2-7B-chat used here, we optimize tokens in addition to a fixed system prompt and instruction. In this setting, we show that MiniPrompt compresses the quote in red to the two blue tokens in the prompt in the top cell. Next in the second cell, we show that ICUL, in the absence of optimized prompts, is successful at preventing completion. Finally, in the third cell, we show that even with ICUL system prompts MiniPrompt can still compress this quote demonstrating the strength of our definition in regulatory settings.
@@ 4.1 The Illusion of Compliance
As data usage regulation advances, there is an emerging motive for organizations and individuals that serve or release models to make it hard to determine that their models have memorized anything. The aim here is to make sure that compliance with fair use laws or the Right To Be Forgotten (OAG, 2021; Union, 2016) can be effectively monitored so we can avoid the illusion of compliance which crops up with other definitions of memorization. Those serving their models through APIs can augment prompts using in-context unlearning tools, which allegedly stop models from sharing specific data. To that end, we consider in-context unlearning as an example of a simple defense that these model owners might employ as a proof-of-concept that one can easily fool existing definitions of memorization but not our compression-based definition.
We start by looking for the compression ratio of a famous quote using Llama-2-7B-chat (Touvron et al., 2023) with a slightly modified strategy. Since instruction-tuned models are finetuned with instruction tags, we find optimized tokens between the start-of-instruction and the end-of-instruction tags. Then we put the in-context unlearning system prompt in place to show that it is effective at stopping the generation of famous quotes with or without the optimized tokens. Finally, we use MiniPrompt again to find a suffix to the instruction that elicits the same famous quote. In Figure 2, we show examples of each of these steps. See Appendix C for further discussion.
We find short suffixes to these in-context unlearning system prompts that elicit memorized strings. Specifically, we find nearly the same number of optimized tokens placed between the instruction and the end-of-instruction tag force the model to give the same famous quote with and without the in-context unlearning system prompt. This consistency in ACR—and therefore the memorization test—matches our intuition that without changing model weights memorized samples are not forgotten. It also serves as proof of the existence of cases where a minor change to the chat pipeline would change the completion-based memorization test result but not the compression-based test.
@@ 4.2 TOFU: Unlearning and Memorization with Author Profiles
As unlearning progresses, we prompt the model to generate answers to the supposedly unlearned questions and record the portion of data that can be completed and compressed. Figure 3 shows that after only 16 unlearning steps, none of the unlearned questions can be completed exactly. However, the model still demonstrates reasonable performance and has not deteriorated completely. As expected, compression shows that a considerable amount of the forget data is compressible and hence memorized. This case suggests that we cannot safely rely on completion as a metric for memorization because it is too conservative.
@@
Рис. 3: Left: Completion vs compression on TOFU data, unlearning Phi-1.5 with gradient ascent. Right: Generation after 20 unlearning steps.
@@ 4.3 Trying to Forget Harry Potter
In their paper on unlearning Harry Potter, Eldan and Russinovich (2023) claim that Llama-2-chat can forget about Harry Potter with several steps of unlearning. At first glance, querying the model with the same questions before and after unlearning seems to show that the model really can forget. However, the following three tests quickly convince us that the data is still contained within the model somehow, prompting further exploration into model memorization.
@@
    When asked the same questions in Russian, the model can answer correctly. We provide examples of such behavior in Appendix D and Lynch et al. (2024) make the same observation.
@@
    With adversarial attacks designed to force affirmative answers without any information about the true answer, we can elicit the correct response—$57\%$ of the Harry Potter related responses can be elicited from the original Llama-2 model, and $50\%$ can still be elicited after unlearning (Figure 7).
@@
Рис. 4: Negative log-likelihood (normalized to $[0,1]$) of true and false answers given a Harry Potter question. Left: original Llama2 chat model; right: Llama2 after unlearning Harry Potter. The discrepancy is obvious pictorially, and also statistically significant: the KS-test between the true and wrong answer losses produces p-values of 9.7e-24 and 5.9e-14, respectively.
@@ 4.4 Bigger Models Memorize More
Since prior work has proposed alternative definitions of memorization that show that bigger models memorize more (Carlini et al., 2023), we ask whether our definition leads to the same finding. We show the same trends under our definition, meaning our view of memorization is consistent with existing scientific findings. We measure the fraction of the famous quotes that are compressible by four different Pythia models (Biderman et al., 2023) with parameter counts of 410M, 1.4B, 6.9B, and 12B and the results are in Figure 5.
@@ 4.5 Validation of MiniPrompt with Four Categories of Data
We look at random sequences of tokens because we want to rule out the possibility that we can always find an adversarial, few-token prompt even for random output—random strings should not be compressible. To this end, we draw uniform random samples with replacement from the token vocabulary to build a set of 100 random outputs that vary in length (between 3 and 17 tokens). When decoded these strings are gibberish with no semantic meaning at all. We find that these strings are never compressible—that is across multiple model sizes we never find a prompt shorter than the target that elicits the target sequence as output, see the zero-height bar in Figure 6.
Рис. 5: Memorization in Pythia models. Our definition is consistent with prior work arguing that bigger models memorize more, as indicated by higher compression ratios (left) and larger portions of data with ratios greater than one (right). These figures are from the Famous Quotes dataset.
Рис. 6: Memorization in Pythia-1.4B. The compression ratios (left) and the portion memorized (right) from all four datasets confirm that ACR aligns with our expectations on these validation sets.
@@
To further determine the validity of our definition, we investigate the average compression rate of natural text that is not in the training set. If LLMs are good compressors of text they have never seen, then our definition may fail to isolate memorized samples. We take random sentences from Associated Press articles that were published in November 2023, well after the models we experiment with were trained. These strings are samples from the distribution of training data as the training set includes real news articles from just a few months prior. Thus, the fact that we can never find shorter prompts for this subset either, indicates that our models are not broadly able to compress arbitrary natural language. Again, see the zero-height bar in Figure 6.
@@
Next, we turn our attention to famous strings, of which many should be categorized as memorized by any useful definition. These are quotes like ‘‘to be, or not to be, that is the question,’’ which are examples repeated many times in the training data. We find that Pythia-1.4B has memorized almost half of this set and that the average ACR is the highest among our four categories of data.
@@
Finally, we look at the memorization of training samples that are not common or famous, but that do exist in the training set. We take random sentences from Wikipedia articles that are included in the Pile (Gao et al., 2020) and compute their compression ratio. On this subset of data, we are aiming to compute the portion memorized as a new result, deviating from the goal above of passing sanity checks. Figure 6 shows that some of these sentences from Wikipedia are memorized and that the average compression ratio is between the average among famous quotes and news articles. Note that the memorized samples form this subset are strings that appear many times on the internet like ‘‘The Burton is a historic apartment building located at Indianapolis, Indiana.’’
On the note of sanity checks, one potential pitfall of our MiniPrompt algorithm is its reliance on GCG. It is possible that there exist shorter strings than we can find. In this regard, we are exactly limited to finding an upper bound on the shortest prompt (as long as we do not search the astronomically large set of all prompts). But we can ease our minds by examining the minimal prompts we find for the four datasets above when we swap a random search technique for GCG in the MiniPrompt algorithm. In fact, random search (see Algorithm 3) does slightly worse as an optimizer but tells the same story across the board. Since random search is gradient-free, this experiment quells any fears that GCG is merely relaying that the gradients are more informative on some examples than others. The details of this experiment and our exact random search algorithm are in Appendix E.
@@ 5 Discussion
Limitations
Our findings are limited in that we mostly consider Pythia models and a natural question we do not address is what kinds of things are memorized by prominent state-of-the-art models. Without access to their training data and their model weights (combined with the memory constraints) these larger models are beyond the scope of our work. Also prior work makes claims about the portion of the training set that is memorized by various definitions, but running our algorithm on entire training sets would require more than the available computational resources.
Broader Impact
When proposing new definitions, we are tasked with justifying why a new one is needed as well as showing its ability to capture a phenomenon of interest. This stands in contrast to developing detection/classification tools whose accuracy can easily be measured using labeled data. It is difficult by nature to define memorization as there is no set of ground truth labels that indicate which samples are memorized. Consequently, the criteria for a memorization definition should rely on how useful it is. Our definition is a promising direction for future regulation on LLM fair use of data as well as helping model owners confidently release models trained on sensitive data without releasing that data. Deploying our framework in practice may require careful thought about how to set the compression threshold but as it relates to the legal setting this is not a limitation as law suits always have some subjectivity (Downing, 2024). Our hope is to provide regulators, model owners, and the courts a mechanism to measure the extent to which a model contains a particular string within its weights and make discussion about data usage more grounded and quantitative.
Acknowledgments
We thank Florian Tramèr, Vaishnavh Nagarajan, and Alvaro Velasquez for their constructive input and creative edge cases.
Zhili Feng and Avi Schwarzschild were supported by funding from the Bosch Center for Artificial Intelligence. Pratyush Maini was supported by DARPA GARD Contract HR00112020006.
@@ Acknowledgments
  • Biderman et al. (2023) Stella Biderman, Hailey Schoelkopf, Quentin Gregory Anthony, Herbie Bradley, Kyle O’Brien, Eric Hallahan, Mohammad Aflah Khan, Shivanshu Purohit, USVSN Sai Prashanth, Edward Raff, et al. Pythia: A suite for analyzing large language models across training and scaling. In International Conference on Machine Learning, pages 2397–2430. PMLR, 2023.
  • Bourtoule et al. (2021) Lucas Bourtoule, Varun Chandrasekaran, Christopher A Choquette-Choo, Hengrui Jia, Adelin Travers, Baiwu Zhang, David Lie, and Nicolas Papernot. Machine unlearning. In 2021 IEEE Symposium on Security and Privacy (SP), pages 141–159. IEEE, 2021.
  • Carlini et al. (2019) Nicholas Carlini, Chang Liu, Úlfar Erlingsson, Jernej Kos, and Dawn Song. The secret sharer: Evaluating and testing unintended memorization in neural networks. In 28th USENIX security symposium (USENIX security 19), pages 267–284, 2019.
@@
  • Choi et al. (2024) Dami Choi, Yonadav Shavit, and David K Duvenaud. Tools for verifying neural models’ training data. Advances in Neural Information Processing Systems, 36, 2024.
@@
  • Downing (2024) Kate Downing. Copyright fundamentals for ai researchers. In Proceedings of the Twelfth International Conference on Learning Representations (ICLR), 2024. URL https://iclr.cc/media/iclr-2024/Slides/21804.pdf.
@@
  • Eldan and Russinovich (2023) Ronen Eldan and Mark Russinovich. Who’s harry potter? approximate unlearning in llms. arXiv preprint arXiv:2310.02238, 2023.
  • Gailly and Adler (1992) Jean-Loup Gailly and Mark Adler. gzip. https://www.gnu.org/software/gzip/, 1992. Accessed: 2024-05-21.
@@
  • Ippolito et al. (2023) Daphne Ippolito, Florian Tramèr, Milad Nasr, Chiyuan Zhang, Matthew Jagielski, Katherine Lee, Christopher A Choquette-Choo, and Nicholas Carlini. Preventing generation of verbatim memorization in language models gives a false sense of privacy. In Proceedings of the 16th International Natural Language Generation Conference, pages 28–53. Association for Computational Linguistics, 2023.
@@
  • Lynch et al. (2024) Aengus Lynch, Phillip Guo, Aidan Ewart, Stephen Casper, and Dylan Hadfield-Menell. Eight methods to evaluate robust unlearning in llms. arXiv preprint arXiv:2402.16835, 2024.
@@
  • Metz and Robertson (2024) Cade Metz and Katie Robertson. Openai seeks to dismiss parts of the new york times’s lawsuit. The New York Times, 2024. URL https://www.nytimes.com/2024/02/27/technology/openai-new-york-times-lawsuit.html.
@@
  • Pham et al. (2024) Minh Pham, Kelly O Marshall, Chinmay Hegde, and Niv Cohen. Robust concept erasure using task vectors. arXiv preprint arXiv:2404.03631, 2024.
  • Sanfilippo (2006) Salvatore Sanfilippo. Smaz: Small strings compression library. https://github.com/antirez/smaz, 2006. Accessed: 2024-05-21.
  • Schwinn et al. (2024) Leo Schwinn, David Dobre, Sophie Xhonneux, Gauthier Gidel, and Stephan Gunnemann. Soft prompt threats: Attacking safety alignment and unlearning in open-source llms through the embedding space. arXiv preprint arXiv:2402.09063, 2024.
@@
  • Shokri et al. (2017) Reza Shokri, Marco Stronati, Congzheng Song, and Vitaly Shmatikov. Membership inference attacks against machine learning models. In 2017 IEEE symposium on security and privacy (SP), pages 3–18. IEEE, 2017.
@@
  • Ullah et al. (2021) Enayat Ullah, Tung Mai, Anup Rao, Ryan A Rossi, and Raman Arora. Machine unlearning via algorithmic stability. In Conference on Learning Theory, pages 4126–4142. PMLR, 2021.
@@
Приложение A Additional Related Work
Prompt Optimization
We borrow prompt optimization tools from work on jailbreaking where the goal is to force LLMs to break their alignment and produce nefarious and toxic output by way of optimizing prompts [Zou et al., 2023, Zhu et al., 2023, Chao et al., 2023, Andriushchenko, 2023]. Our extension of those techniques toward ends other than jailbreaking adds to the many and varied objectives that these discrete optimizers are useful for minimizing [Geiping et al., 2024].
Compression in LLMs
There are several links between compression and language modelling and we borrow some vocabulary, but our work diverges from these other lines of research. For example, Delétang et al. [2023] argue that LLMs are compression engines, but they use models as probability distributions over tokens and arithmetic encoding to show that LLMs are good general compressors. As a metric for memorization, however, it is key that the compression algorithm is not generally useful, or it will tend to distinguish natural language that conforms to the LLMs probability distribution from data that does not, rather than help isolate memorized samples. Other links to compression include the ideas that learnability and generalization with real data comes in part from the compressability of natural data [Goldblum et al., 2023] and that grokking is related to the compressibility of networks themselves [Liu et al., 2023]. Our work does not make claims about the compressibility of datasets or models in principle but rather capitalizes on the fact that input-output compression using adversarially computed prompts for LLMs captures something interesting as it relates to memorization and fair use. In fact, Jiang et al. [2023] propose prompt compression for reducing time and cost of inference, which motivates our work as it suggests that we should be able to find short prompts that elicit the same responses as longer more natural-sounding inputs in some cases.
Unlearning
The focus of machine unlearning [Bourtoule et al., 2021, Sekhari et al., 2021, Ullah et al., 2021, Pham et al., 2024, Lynch et al., 2024, Schwinn et al., 2024] is to remove private, sensitive, or false data from models without retraining them from scratch. Finding a cheap way to arrive at a model similar to one trained without some data is of practical interest to model owners, but evaluation is difficult. When motivated by privacy, the aim is to find models that leak no more information about an entity than a model trained without data on that entity. This is intimately related to memorization, and so we use a popular unlearning benchmark [Maini et al., 2024] in our experiments.
Приложение B Algorithms In Our Experiments
@@
Algorithm 2 Greedy Coordinate Gradient (GCG) [Zou et al., 2023]
@@
Algorithm 3 Random Search (for LLM prompts) [Andriushchenko, 2023]
@@
Приложение C More Details of In-context Unlearning
@@
In the textbox below, note that the ICUL system prompt leads to an abstention response with the original optimized prompt.
[…]
Response: I apologize, but I cannot provide you with a quote that promotes…
Приложение D More Details of Unlearning Harry Potter
[…]
Generation: "It is our choices, Harry, that show what we truly are, far more than our abil%ities." - Albus Dumbledore
Рис. 7: Left: Fraction of Harry Potter texts that are compressible. Right: an example of hard tokens that elicit Harry Potter text.